On a summer day the pressure is 1 atm, the temperature is 32°C, and the relative humidity is 95 percent. This air is to be conditioned to 24°C and 60 percent relative humidity.

a) Determine the amount of cooling, in kJ, required and

b) Water removed, in kg, per 1000 m3 of dry air processed at the entrance to the system. Cooling coils Ti 32С ф,-95% т,-24°C φ2 60% l atm Condensate
Answer #1

Detailed solution is shown below ask if any doubt

Frem Poycometitc chas », = o.qo5 m3l 2. 2. Pom satwrated tables at T28 28-25 30-2s hw = 117-316 KJ/kg.mw19.668 K oul out \104.97 ((06.8-52.1)-19.66 8(n7-376) = 59778-817一々308-55 out- ebut = 574-70-327 kJ